The direction cosines of the line joining the points A(-6,5,-4),B(-5,2,-4) is ?

Grade:12

1 Answers

Nishant Vora IIT Patna
7 years ago
DR are given by\vec{AB}= (1,-3,0)

hence DC are given by\hat{AB}=(-1 / \sqrt{10}, 3/ \sqrt{10}, 0)

Thanks